OpenStudy (anonymous):

Write and equation for the line in slope-intercept form. Passing through (5,2) and parallel to the line whose equation is y=-7x+6.

OpenStudy (anonymous):

When you have an equation of a line in the form of \(y=mx+c\), you can take the slope as being the number in front of the \(x\) term. Then you can use the equation of a line formula: \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

I know the answer is y=-7x+37, I'm just trying to figure out how to get there.

OpenStudy (anonymous):

well use the equation of a line formula dalvoron gave you, along with the information provided in the problem. In order to have a line that is parallel they must have the same slope, in this case the slope is m=-7 and we have the points (5,2) (x1,y1) now you just plug in and solve for y

OpenStudy (anonymous):

Thank you, you beautiful man.

OpenStudy (anonymous):

So we are already given the slope from the parallel line \[m=-7\] We also know that the equation of a line is \[y=mx+b\]And we know an (x,y) coordinate (5,2) We can plug in the coordinate point and the slow into the line equation and solve for b:\[y=mx+b\]\[b=y-mx\]Plug in x, y, and m:\[b=5-(-7)(2)\]\[b=19\] Thus\[y=-7x+19\]

OpenStudy (anonymous):

y-2=-7(x-5) y-2=-7x+35 y=-7x+37

OpenStudy (anonymous):

y-2=-7(x-5) y-2=-7x+35 y=-7x+37

OpenStudy (anonymous):

b=2-(-7)(5)